\begin{thebibliography}{59}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Anastasopoulos and Chiang(2018)}]{anastasopoulos2018tied}
Antonios Anastasopoulos and David Chiang. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1008} {Tied multitask learning
  for neural speech translation}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 82--91, New Orleans, Louisiana.
  Association for Computational Linguistics.

\bibitem[{Ansari et~al.(2020)Ansari, Axelrod, Bach, Bojar, Cattoni, Dalvi,
  Durrani, Federico, Federmann, Gu, Huang, Knight, Ma, Nagesh, Negri, Niehues,
  Pino, Salesky, Shi, St{\"u}ker, Turchi, Waibel, and
  Wang}]{ansari-etal-2020-findings}
Ebrahim Ansari, Amittai Axelrod, Nguyen Bach, Ond{\v{r}}ej Bojar, Roldano
  Cattoni, Fahim Dalvi, Nadir Durrani, Marcello Federico, Christian Federmann,
  Jiatao Gu, Fei Huang, Kevin Knight, Xutai Ma, Ajay Nagesh, Matteo Negri, Jan
  Niehues, Juan Pino, Elizabeth Salesky, Xing Shi, Sebastian St{\"u}ker, Marco
  Turchi, Alexander Waibel, and Changhan Wang. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.iwslt-1.1} {{FINDINGS} {OF}
  {THE} {IWSLT} 2020 {EVALUATION} {CAMPAIGN}}.
\newblock In \emph{Proceedings of the 17th International Conference on Spoken
  Language Translation}, pages 1--34, Online. Association for Computational
  Linguistics.

\bibitem[{Arivazhagan et~al.(2019{\natexlab{a}})Arivazhagan, Bapna, Firat,
  Lepikhin, Johnson, Krikun, Chen, Cao, Foster, Cherry
  et~al.}]{arivazhagan2019massively}
Naveen Arivazhagan, Ankur Bapna, Orhan Firat, Dmitry Lepikhin, Melvin Johnson,
  Maxim Krikun, Mia~Xu Chen, Yuan Cao, George Foster, Colin Cherry, et~al.
  2019{\natexlab{a}}.
\newblock Massively multilingual neural machine translation in the wild:
  Findings and challenges.
\newblock \emph{arXiv preprint arXiv:1907.05019}.

\bibitem[{Arivazhagan et~al.(2019{\natexlab{b}})Arivazhagan, Cherry, Macherey,
  Chiu, Yavuz, Pang, Li, and Raffel}]{arivazhagan2019monotonic}
Naveen Arivazhagan, Colin Cherry, Wolfgang Macherey, Chung-Cheng Chiu, Semih
  Yavuz, Ruoming Pang, Wei Li, and Colin Raffel. 2019{\natexlab{b}}.
\newblock Monotonic infinite lookback attention for simultaneous machine
  translation.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 1313--1323.

\bibitem[{Baevski et~al.(2020)Baevski, Zhou, Mohamed, and
  Auli}]{NEURIPS2020_92d1e1eb}
Alexei Baevski, Yuhao Zhou, Abdelrahman Mohamed, and Michael Auli. 2020.
\newblock \href
  {https://proceedings.neurips.cc/paper/2020/file/92d1e1eb1cd6f9fba3227870bb6d7f07-Paper.pdf}
  {wav2vec 2.0: A framework for self-supervised learning of speech
  representations}.
\newblock In \emph{Advances in Neural Information Processing Systems},
  volume~33, pages 12449--12460. Curran Associates, Inc.

\bibitem[{Bansal et~al.(2019)Bansal, Kamper, Livescu, Lopez, and
  Goldwater}]{bansal2018pre}
Sameer Bansal, Herman Kamper, Karen Livescu, Adam Lopez, and Sharon Goldwater.
  2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1006} {Pre-training on
  high-resource speech recognition improves low-resource speech-to-text
  translation}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 58--68, Minneapolis,
  Minnesota. Association for Computational Linguistics.

\bibitem[{B{\'e}rard et~al.(2018)B{\'e}rard, Besacier, Kocabiyikoglu, and
  Pietquin}]{berard2018end}
Alexandre B{\'e}rard, Laurent Besacier, Ali~Can Kocabiyikoglu, and Olivier
  Pietquin. 2018.
\newblock End-to-end automatic speech translation of audiobooks.
\newblock In \emph{2018 IEEE International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 6224--6228. IEEE.

\bibitem[{B{\'e}rard et~al.(2016)B{\'e}rard, Pietquin, Servan, and
  Besacier}]{berard2016listen}
Alexandre B{\'e}rard, Olivier Pietquin, Christophe Servan, and Laurent
  Besacier. 2016.
\newblock Listen and translate: A proof of concept for end-to-end
  speech-to-text translation.
\newblock \emph{arXiv preprint arXiv:1612.01744}.

\bibitem[{Can et~al.(2018)Can, Martinez, Papadopoulos, and Narayanan}]{pykaldi}
Dogan Can, Victor~R. Martinez, Pavlos Papadopoulos, and Shrikanth~S. Narayanan.
  2018.
\newblock Pykaldi: A python wrapper for kaldi.
\newblock In \emph{Acoustics, Speech and Signal Processing (ICASSP), 2018 IEEE
  International Conference on}. IEEE.

\bibitem[{Chan et~al.(2016)Chan, Jaitly, Le, and Vinyals}]{chan2016listen}
William Chan, Navdeep Jaitly, Quoc Le, and Oriol Vinyals. 2016.
\newblock Listen, attend and spell: A neural network for large vocabulary
  conversational speech recognition.
\newblock In \emph{2016 IEEE International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 4960--4964. IEEE.

\bibitem[{Cherry and Foster(2019)}]{cherry2019thinking}
Colin Cherry and George Foster. 2019.
\newblock Thinking slow about latency evaluation for simultaneous machine
  translation.
\newblock \emph{arXiv preprint arXiv:1906.00048}.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'a}n, Grave, Ott, Zettlemoyer, and Stoyanov}]{conneau2019unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'a}n, Edouard Grave, Myle Ott, Luke Zettlemoyer, and
  Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.747} {Unsupervised
  cross-lingual representation learning at scale}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 8440--8451, Online. Association for
  Computational Linguistics.

\bibitem[{{Di Gangi} et~al.(2019){Di Gangi}, {Negri}, and {Turchi}}]{9004003}
M.~A. {Di Gangi}, M.~{Negri}, and M.~{Turchi}. 2019.
\newblock One-to-many multilingual end-to-end speech translation.
\newblock In \emph{2019 IEEE Automatic Speech Recognition and Understanding
  Workshop (ASRU)}, pages 585--592.

\bibitem[{Di~Gangi et~al.(2019{\natexlab{a}})Di~Gangi, Cattoni, Bentivogli,
  Negri, and Turchi}]{di2019must}
Mattia~A Di~Gangi, Roldano Cattoni, Luisa Bentivogli, Matteo Negri, and Marco
  Turchi. 2019{\natexlab{a}}.
\newblock Must-c: a multilingual speech translation corpus.
\newblock In \emph{2019 Conference of the North American Chapter of the
  Association for Computational Linguistics: Human Language Technologies},
  pages 2012--2017. Association for Computational Linguistics.

\bibitem[{Di~Gangi et~al.(2019{\natexlab{b}})Di~Gangi, Negri, Cattoni, Dessi,
  and Turchi}]{di-gangi-etal-2019-enhancing}
Mattia~Antonino Di~Gangi, Matteo Negri, Roldano Cattoni, Roberto Dessi, and
  Marco Turchi. 2019{\natexlab{b}}.
\newblock \href {https://www.aclweb.org/anthology/W19-6603} {Enhancing
  transformer for end-to-end speech-to-text translation}.
\newblock In \emph{Proceedings of Machine Translation Summit XVII Volume 1:
  Research Track}, pages 21--31, Dublin, Ireland. European Association for
  Machine Translation.

\bibitem[{Duong et~al.(2016)Duong, Anastasopoulos, Chiang, Bird, and
  Cohn}]{duong2016attentional}
Long Duong, Antonios Anastasopoulos, David Chiang, Steven Bird, and Trevor
  Cohn. 2016.
\newblock An attentional model for speech translation without transcription.
\newblock In \emph{Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 949--959.

\bibitem[{Graves et~al.(2006)Graves, Fern{\'a}ndez, Gomez, and
  Schmidhuber}]{graves2006connectionist}
Alex Graves, Santiago Fern{\'a}ndez, Faustino Gomez, and J{\"u}rgen
  Schmidhuber. 2006.
\newblock Connectionist temporal classification: labelling unsegmented sequence
  data with recurrent neural networks.
\newblock In \emph{Proceedings of the 23rd international conference on Machine
  learning}, pages 369--376.

\bibitem[{Gulati et~al.(2020)Gulati, Qin, Chiu, Parmar, Zhang, Yu, Han, Wang,
  Zhang, Wu, and Pang}]{gulati20_interspeech}
Anmol Gulati, James Qin, Chung-Cheng Chiu, Niki Parmar, Yu~Zhang, Jiahui Yu,
  Wei Han, Shibo Wang, Zhengdong Zhang, Yonghui Wu, and Ruoming Pang. 2020.
\newblock \href {https://doi.org/10.21437/Interspeech.2020-3015} {{Conformer:
  Convolution-augmented Transformer for Speech Recognition}}.
\newblock In \emph{Proc. Interspeech 2020}, pages 5036--5040.

\bibitem[{Harris et~al.(2020)Harris, Millman, van~der Walt, Gommers, Virtanen,
  Cournapeau, Wieser, Taylor, Berg, Smith et~al.}]{harris2020array}
Charles~R Harris, K~Jarrod Millman, St{\'e}fan~J van~der Walt, Ralf Gommers,
  Pauli Virtanen, David Cournapeau, Eric Wieser, Julian Taylor, Sebastian Berg,
  Nathaniel~J Smith, et~al. 2020.
\newblock Array programming with numpy.
\newblock \emph{Nature}, 585(7825):357--362.

\bibitem[{Inaguma et~al.(2020)Inaguma, Kiyono, Duh, Karita, Yalta, Hayashi, and
  Watanabe}]{inaguma-etal-2020-espnet}
Hirofumi Inaguma, Shun Kiyono, Kevin Duh, Shigeki Karita, Nelson Yalta, Tomoki
  Hayashi, and Shinji Watanabe. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.acl-demos.34}
  {{ESP}net-{ST}: All-in-one speech translation toolkit}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics: System Demonstrations}, pages 302--311,
  Online. Association for Computational Linguistics.

\bibitem[{Kahn et~al.(2020)Kahn, Lee, and Hannun}]{Kahn2020SelfTrainingFE}
Jacob Kahn, Ann Lee, and Awni Hannun. 2020.
\newblock Self-training for end-to-end speech recognition.
\newblock \emph{ICASSP 2020 - 2020 IEEE International Conference on Acoustics,
  Speech and Signal Processing (ICASSP)}, pages 7084--7088.

\bibitem[{Klein et~al.(2017)Klein, Kim, Deng, Senellart, and
  Rush}]{klein2017opennmt}
Guillaume Klein, Yoon Kim, Yuntian Deng, Jean Senellart, and Alexander Rush.
  2017.
\newblock \href {https://www.aclweb.org/anthology/P17-4012} {{O}pen{NMT}:
  Open-source toolkit for neural machine translation}.
\newblock In \emph{Proceedings of {ACL} 2017, System Demonstrations}, pages
  67--72, Vancouver, Canada. Association for Computational Linguistics.

\bibitem[{Ko et~al.(2017)Ko, Peddinti, Povey, Seltzer, and
  Khudanpur}]{ko2017study}
Tom Ko, Vijayaditya Peddinti, Daniel Povey, Michael~L Seltzer, and Sanjeev
  Khudanpur. 2017.
\newblock A study on data augmentation of reverberant speech for robust speech
  recognition.
\newblock In \emph{2017 IEEE International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 5220--5224. IEEE.

\bibitem[{Kuchaiev et~al.(2018)Kuchaiev, Ginsburg, Gitman, Lavrukhin, Case, and
  Micikevicius}]{kuchaiev2018openseq2seq}
Oleksii Kuchaiev, Boris Ginsburg, Igor Gitman, Vitaly Lavrukhin, Carl Case, and
  Paulius Micikevicius. 2018.
\newblock Openseq2seq: extensible toolkit for distributed and mixed precision
  training of sequence-to-sequence models.
\newblock In \emph{Proceedings of Workshop for NLP Open Source Software
  (NLP-OSS)}, pages 41--46.

\bibitem[{Kudo and Richardson(2018)}]{kudo-richardson-2018-sentencepiece}
Taku Kudo and John Richardson. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-2012} {{S}entence{P}iece: A
  simple and language independent subword tokenizer and detokenizer for neural
  text processing}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing: System Demonstrations}, pages 66--71, Brussels,
  Belgium. Association for Computational Linguistics.

\bibitem[{Lewis et~al.(2020)Lewis, Liu, Goyal, Ghazvininejad, Mohamed, Levy,
  Stoyanov, and Zettlemoyer}]{lewis2019bart}
Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed,
  Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.703} {{BART}:
  Denoising sequence-to-sequence pre-training for natural language generation,
  translation, and comprehension}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 7871--7880, Online. Association for
  Computational Linguistics.

\bibitem[{Li et~al.(2019)Li, Zhang, Sainath, Wu, and Chan}]{li2019bytes}
Bo~Li, Yu~Zhang, Tara Sainath, Yonghui Wu, and William Chan. 2019.
\newblock Bytes are all you need: End-to-end multilingual speech recognition
  and synthesis with bytes.
\newblock In \emph{ICASSP 2019-2019 IEEE International Conference on Acoustics,
  Speech and Signal Processing (ICASSP)}, pages 5621--5625. IEEE.

\bibitem[{Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov}]{liu2019roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019.
\newblock Roberta: A robustly optimized bert pretraining approach.
\newblock \emph{arXiv preprint arXiv:1907.11692}.

\bibitem[{Liu et~al.(2020)Liu, Zhang, Xiong, Zhou, He, Wu, Wang, and
  Zong}]{liu2019synchronous}
Yuchen Liu, Jiajun Zhang, Hao Xiong, Long Zhou, Zhongjun He, Hua Wu, Haifeng
  Wang, and Chengqing Zong. 2020.
\newblock \href {https://doi.org/10.1609/aaai.v34i05.6360} {Synchronous speech
  recognition and speech-to-text translation with interactive decoding}.
\newblock \emph{Proceedings of the AAAI Conference on Artificial Intelligence},
  34:8417--8424.

\bibitem[{Ma et~al.(2019)Ma, Huang, Xiong, Zheng, Liu, Zheng, Zhang, He, Liu,
  Li, Wu, and Wang}]{Ma2018STACLST}
Mingbo Ma, Liang Huang, Hao Xiong, Renjie Zheng, Kaibo Liu, Baigong Zheng,
  Chuanqiang Zhang, Zhongjun He, Hairong Liu, Xing Li, Hua Wu, and Haifeng
  Wang. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1289} {{STACL}: Simultaneous
  translation with implicit anticipation and controllable latency using
  prefix-to-prefix framework}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 3025--3036, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Ma et~al.(2020{\natexlab{a}})Ma, Dousti, Wang, Gu, and
  Pino}]{ma2020simuleval}
Xutai Ma, Mohammad~Javad Dousti, Changhan Wang, Jiatao Gu, and Juan Pino.
  2020{\natexlab{a}}.
\newblock Simuleval: An evaluation toolkit for simultaneous translation.
\newblock \emph{arXiv preprint arXiv:2007.16193}.

\bibitem[{Ma et~al.(2020{\natexlab{b}})Ma, Pino, Cross, Puzon, and
  Gu}]{ma2019monotonic}
Xutai Ma, Juan Pino, James Cross, Liezl Puzon, and Jiatao Gu.
  2020{\natexlab{b}}.
\newblock Monotonic multihead attention.
\newblock \emph{8th International Conference on Learning Representations,
  {ICLR} 2020, Addis Ababa, Ethiopia, April 26-30, 2020, Conference Track
  Proceedings}.

\bibitem[{Micikevicius et~al.(2018)Micikevicius, Narang, Alben, Diamos, Elsen,
  Garcia, Ginsburg, Houston, Kuchaiev, Venkatesh
  et~al.}]{micikevicius2018mixed}
Paulius Micikevicius, Sharan Narang, Jonah Alben, Gregory Diamos, Erich Elsen,
  David Garcia, Boris Ginsburg, Michael Houston, Oleksii Kuchaiev, Ganesh
  Venkatesh, et~al. 2018.
\newblock Mixed precision training.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Mohamed et~al.(2019)Mohamed, Okhonko, and
  Zettlemoyer}]{Mohamed2019TransformersWC}
Abdelrahman Mohamed, Dmytro Okhonko, and Luke Zettlemoyer. 2019.
\newblock Transformers with convolutional context for asr.
\newblock \emph{arXiv preprint arXiv:1904.11660}.

\bibitem[{Ng et~al.(2019)Ng, Yee, Baevski, Ott, Auli, and
  Edunov}]{ng-etal-2019-facebook}
Nathan Ng, Kyra Yee, Alexei Baevski, Myle Ott, Michael Auli, and Sergey Edunov.
  2019.
\newblock \href {https://doi.org/10.18653/v1/W19-5333} {{F}acebook {FAIR}{'}s
  {WMT}19 news translation task submission}.
\newblock In \emph{Proceedings of the Fourth Conference on Machine Translation
  (Volume 2: Shared Task Papers, Day 1)}, pages 314--319, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Ott et~al.(2019)Ott, Edunov, Baevski, Fan, Gross, Ng, Grangier, and
  Auli}]{ott2019fairseq}
Myle Ott, Sergey Edunov, Alexei Baevski, Angela Fan, Sam Gross, Nathan Ng,
  David Grangier, and Michael Auli. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-4009} {fairseq: A fast,
  extensible toolkit for sequence modeling}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics (Demonstrations)},
  pages 48--53, Minneapolis, Minnesota. Association for Computational
  Linguistics.

\bibitem[{Panayotov et~al.(2015)Panayotov, Chen, Povey, and
  Khudanpur}]{panayotov2015librispeech}
Vassil Panayotov, Guoguo Chen, Daniel Povey, and Sanjeev Khudanpur. 2015.
\newblock Librispeech: an asr corpus based on public domain audio books.
\newblock In \emph{2015 IEEE International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 5206--5210. IEEE.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of the 40th annual meeting of the Association
  for Computational Linguistics}, pages 311--318.

\bibitem[{Park et~al.(2019)Park, Chan, Zhang, Chiu, Zoph, Cubuk, and
  Le}]{park2019specaugment}
Daniel~S. Park, William Chan, Yu~Zhang, Chung-Cheng Chiu, Barret Zoph, Ekin~D.
  Cubuk, and Quoc~V. Le. 2019.
\newblock \href {https://doi.org/10.21437/interspeech.2019-2680} {Specaugment:
  A simple data augmentation method for automatic speech recognition}.
\newblock \emph{Interspeech 2019}.

\bibitem[{Paszke et~al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan,
  Killeen, Lin, Gimelshein, Antiga et~al.}]{paszke2019pytorch}
Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory
  Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, et~al.
  2019.
\newblock Pytorch: An imperative style, high-performance deep learning library.
\newblock In \emph{Advances in neural information processing systems}, pages
  8026--8037.

\bibitem[{Pino et~al.(2020)Pino, Xu, Ma, Dousti, and
  Tang}]{pino2020selftraining}
Juan Pino, Qiantong Xu, Xutai Ma, Mohammad~Javad Dousti, and Yun Tang. 2020.
\newblock Self-training for end-to-end speech translation.
\newblock \emph{arXiv preprint arXiv:2006.02490}.

\bibitem[{Popovi{\'c}(2015)}]{popovic2015chrf}
Maja Popovi{\'c}. 2015.
\newblock chrf: character n-gram f-score for automatic mt evaluation.
\newblock In \emph{Proceedings of the Tenth Workshop on Statistical Machine
  Translation}, pages 392--395.

\bibitem[{Post(2018)}]{post-2018-call}
Matt Post. 2018.
\newblock \href {https://www.aclweb.org/anthology/W18-6319} {A call for clarity
  in reporting {BLEU} scores}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 186--191, Belgium, Brussels. Association for
  Computational Linguistics.

\bibitem[{Povey et~al.(2011)Povey, Ghoshal, Boulianne, Burget, Glembek, Goel,
  Hannemann, Motlicek, Qian, Schwarz et~al.}]{povey2011kaldi}
Daniel Povey, Arnab Ghoshal, Gilles Boulianne, Lukas Burget, Ondrej Glembek,
  Nagendra Goel, Mirko Hannemann, Petr Motlicek, Yanmin Qian, Petr Schwarz,
  et~al. 2011.
\newblock The kaldi speech recognition toolkit.
\newblock In \emph{IEEE 2011 workshop on automatic speech recognition and
  understanding}, CONF. IEEE Signal Processing Society.

\bibitem[{Pratap et~al.(2018)Pratap, Hannun, Xu, Cai, Kahn, Synnaeve,
  Liptchinsky, and Collobert}]{pratap2018w2l}
Vineel Pratap, Awni Hannun, Qiantong Xu, Jeff Cai, Jacob Kahn, Gabriel
  Synnaeve, Vitaliy Liptchinsky, and Ronan Collobert. 2018.
\newblock \href {https://arxiv.org/abs/1812.07625} {wav2letter++: The fastest
  open-source speech recognition system}.
\newblock \emph{CoRR}, abs/1812.07625.

\bibitem[{Raffel et~al.(2017)Raffel, Luong, Liu, Weiss, and
  Eck}]{raffel2017online}
Colin Raffel, Minh-Thang Luong, Peter~J Liu, Ron~J Weiss, and Douglas Eck.
  2017.
\newblock Online and linear-time attention by enforcing monotonic alignments.
\newblock In \emph{International Conference on Machine Learning}, pages
  2837--2846.

\bibitem[{Schneider et~al.(2019)Schneider, Baevski, Collobert, and
  Auli}]{Schneider2019}
Steffen Schneider, Alexei Baevski, Ronan Collobert, and Michael Auli. 2019.
\newblock \href {https://doi.org/10.21437/Interspeech.2019-1873} {{wav2vec:
  Unsupervised Pre-Training for Speech Recognition}}.
\newblock In \emph{Proc. Interspeech 2019}, pages 3465--3469.

\bibitem[{Shen et~al.(2019)Shen, Nguyen, Wu, Chen, Chen, Jia, Kannan, Sainath,
  Cao, Chiu et~al.}]{shen2019lingvo}
Jonathan Shen, Patrick Nguyen, Yonghui Wu, Zhifeng Chen, Mia~X Chen, Ye~Jia,
  Anjuli Kannan, Tara Sainath, Yuan Cao, Chung-Cheng Chiu, et~al. 2019.
\newblock Lingvo: a modular and scalable framework for sequence-to-sequence
  modeling.
\newblock \emph{arXiv preprint arXiv:1902.08295}.

\bibitem[{Synnaeve et~al.(2019)Synnaeve, Xu, Kahn, Grave, Likhomanenko, Pratap,
  Sriram, Liptchinsky, and Collobert}]{synnaeve2019end}
Gabriel Synnaeve, Qiantong Xu, Jacob Kahn, Edouard Grave, Tatiana Likhomanenko,
  Vineel Pratap, Anuroop Sriram, Vitaliy Liptchinsky, and Ronan Collobert.
  2019.
\newblock End-to-end asr: from supervised to semi-supervised learning with
  modern architectures.
\newblock \emph{arXiv preprint arXiv:1911.08460}.

\bibitem[{Vaswani et~al.(2018)Vaswani, Bengio, Brevdo, Chollet, Gomez, Gouws,
  Jones, Kaiser, Kalchbrenner, Parmar, Sepassi, Shazeer, and
  Uszkoreit}]{vaswani2018tensor2tensor}
Ashish Vaswani, Samy Bengio, Eugene Brevdo, Francois Chollet, Aidan Gomez,
  Stephan Gouws, Llion Jones, {\L}ukasz Kaiser, Nal Kalchbrenner, Niki Parmar,
  Ryan Sepassi, Noam Shazeer, and Jakob Uszkoreit. 2018.
\newblock \href {https://www.aclweb.org/anthology/W18-1819} {{T}ensor2{T}ensor
  for neural machine translation}.
\newblock In \emph{Proceedings of the 13th Conference of the Association for
  Machine Translation in the {A}mericas (Volume 1: Research Papers)}, pages
  193--199, Boston, MA. Association for Machine Translation in the Americas.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock In \emph{Advances in neural information processing systems}, pages
  5998--6008.

\bibitem[{Wang et~al.(2020{\natexlab{a}})Wang, Cho, and Gu}]{wang2020aaai}
Changhan Wang, Kyunghyun Cho, and Jiatao Gu. 2020{\natexlab{a}}.
\newblock \href {https://doi.org/10.1609/aaai.v34i05.6451} {Neural machine
  translation with byte-level subwords}.
\newblock \emph{Proceedings of the AAAI Conference on Artificial Intelligence},
  34:9154--9160.

\bibitem[{Wang et~al.(2019)Wang, Jain, Chen, and Gu}]{wang-etal-2019-vizseq}
Changhan Wang, Anirudh Jain, Danlu Chen, and Jiatao Gu. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-3043} {{V}iz{S}eq: a visual
  analysis toolkit for text generation tasks}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP): System Demonstrations}, pages
  253--258, Hong Kong, China. Association for Computational Linguistics.

\bibitem[{Wang et~al.(2020{\natexlab{b}})Wang, Pino, and
  Gu}]{wang2020improving}
Changhan Wang, Juan Pino, and Jiatao Gu. 2020{\natexlab{b}}.
\newblock Improving cross-lingual transfer learning for end-to-end speech
  recognition with speech translation.
\newblock \emph{arXiv preprint arXiv:2006.05474}.

\bibitem[{Wang et~al.(2020{\natexlab{c}})Wang, Wu, and Pino}]{wang2020covost}
Changhan Wang, Anne Wu, and Juan Pino. 2020{\natexlab{c}}.
\newblock Covost 2 and massively multilingual speech-to-text translation.
\newblock \emph{arXiv e-prints}, pages arXiv--2007.

\bibitem[{Wu et~al.(2020)Wu, Wang, Pino, and Gu}]{wu2020selfsupervised}
Anne Wu, Changhan Wang, Juan Pino, and Jiatao Gu. 2020.
\newblock Self-supervised representations improve end-to-end speech
  translation.
\newblock \emph{arXiv preprint arXiv:2006.12124}.

\bibitem[{Zenkel et~al.(2018)Zenkel, Sperber, Niehues, M{\"u}ller, Pham,
  St{\"u}ker, and Waibel}]{zenkel2018open}
Thomas Zenkel, Matthias Sperber, Jan Niehues, Markus M{\"u}ller, Ngoc-Quan
  Pham, Sebastian St{\"u}ker, and Alex Waibel. 2018.
\newblock Open source toolkit for speech to text translation.
\newblock \emph{The Prague Bulletin of Mathematical Linguistics},
  111(1):125--135.

\bibitem[{Zeyer et~al.(2018)Zeyer, Alkhouli, and Ney}]{zeyer-etal-2018-returnn}
Albert Zeyer, Tamer Alkhouli, and Hermann Ney. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-4022} {{RETURNN} as a generic
  flexible neural toolkit with application to translation and speech
  recognition}.
\newblock In \emph{Proceedings of {ACL} 2018, System Demonstrations}, pages
  128--133, Melbourne, Australia. Association for Computational Linguistics.

\bibitem[{Zhang et~al.(2020)Zhang, Qin, Park, Han, Chiu, Pang, Le, and
  Wu}]{zhang2020pushing}
Yu~Zhang, James Qin, Daniel~S Park, Wei Han, Chung-Cheng Chiu, Ruoming Pang,
  Quoc~V Le, and Yonghui Wu. 2020.
\newblock Pushing the limits of semi-supervised learning for automatic speech
  recognition.
\newblock \emph{arXiv preprint arXiv:2010.10504}.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{AACL-IJCNLP2020}
\usepackage{times}
\usepackage{latexsym}

\usepackage{microtype}

\aclfinalcopy % Uncomment this line for the final submission
\usepackage{booktabs}
\usepackage{amssymb}
\usepackage{pifont}
\usepackage{multirow}
\usepackage{graphicx}

\usepackage{color}

\newcommand{\fairseq}{\textsc{fairseq}}
\newcommand{\stot}{\textsc{fairseq S2T}}

\title{\stot: Fast Speech-to-Text Modeling with \fairseq}

\author{Changhan Wang$^1$, Yun Tang$^1$, Xutai Ma$^{1,2}$, Anne Wu$^1$, Sravya Popuri$^1$, \\\textbf{Dmytro Okhonko$^1$, Juan Pino$^1$} \vspace*{0.2cm} \\
$^1$Meta - Fundamental AI Research (FAIR) \\
$^2$Johns Hopkins University \vspace*{0.2cm} \\
  {\tt xutai\_\thinspace ma@jhu.edu} \\
\texttt{\{changhan,yuntang,annewu,spopuri,oxo,juancarabina\}@fb.com} \\
}

\date{}

\begin{document}
\maketitle
\begin{abstract}
We introduce \stot, a \fairseq~\citep{ott2019fairseq} extension for speech-to-text (S2T) modeling tasks such as end-to-end speech recognition and speech-to-text translation. It follows \fairseq's careful design for scalability and extensibility. We provide end-to-end workflows from data pre-processing, model training to offline (online) inference. We implement state-of-the-art RNN-based, Transformer-based as well as Conformer-based models and open-source detailed training recipes. \fairseq's machine translation models and language models can be seamlessly integrated into S2T workflows for multi-task learning or transfer learning. \stot~documentation and examples are available at \url{https://github.com/pytorch/fairseq/tree/master/examples/speech_to_text}.
\end{abstract}

\section{Introduction}

End-to-end sequence-to-sequence (S2S) modeling has witnessed rapidly increased applications in speech-to-text (S2T) tasks. It achieves state-of-the-art performance on automatic speech recognition (ASR)~\citep{park2019specaugment,synnaeve2019end} and leads to the recent resurgence of speech-to-text translation (ST) research~\citep{duong2016attentional,berard2016listen}. ASR and ST are closely related. There are recent attempts to combine the two tasks under the same S2S model architecture via multi-task learning~\citep{anastasopoulos2018tied,liu2019synchronous}. They also benefit from each other via transfer learning~\citep{bansal2018pre,wang2020improving} and are able to leverage additional supervision from machine translation (MT) and language modeling (LM). When supervised data is not abundant, self-supervised pre-training~\citep{Schneider2019,wu2020selfsupervised} and semi-supervised training~\citep{Kahn2020SelfTrainingFE,pino2020selftraining} lowers the requirements on supervision and improves model performance.

\begin{table*}[ht]
    \centering
    \small
    \begin{tabular}{c|ccccccc|cc}
    \toprule
     & \multirow{2}{*}{ASR} & \multirow{2}{*}{LM} & \multirow{2}{*}{MT} & Non-Autoreg. & Offline & Online & Speech & Multi-node & Pre-trained  \\
     & & & & MT & ST & ST & Pre-training & training & models \\
    \midrule 
    ESPNet-ST & \checkmark & \checkmark & \checkmark & & \checkmark & & & \checkmark$^\dagger$ & \checkmark \\
    Lingvo & \checkmark & \checkmark & \checkmark & & \checkmark$^\ddagger$ & & & \checkmark & \\
    OpenSeq2seq$^1$ & \checkmark & \checkmark & \checkmark & & & & & \checkmark & \checkmark \\
    RETURNN$^2$ & \checkmark & \checkmark & \checkmark & & \checkmark & & & \checkmark & \checkmark \\
    SLT.KIT$^3$ & \checkmark & & \checkmark & & \checkmark & & & & \checkmark \\
    Tensor2Tensor$^4$ & \checkmark & \checkmark & \checkmark & & & & & \checkmark & \checkmark \\
    OpenNMT$^5$ & \checkmark & \checkmark & \checkmark & & & & & \checkmark & \checkmark \\
    Kaldi$^6$ & \checkmark & \checkmark & & & & & & & \checkmark \\
    Wav2letter++$^7$ & \checkmark & \checkmark & & & & & & & \checkmark \\
    \midrule
    \textbf{fairseq S2T} & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark \\
    \bottomrule
    \end{tabular}
\caption{Comparison of \stot~with counterpart toolkits (as of July 2020). $^\dagger$ Only available in version 2 (under development). $^\ddagger$ Not publicly available. $^1$ \citet{kuchaiev2018openseq2seq}. $^2$ \citet{zeyer-etal-2018-returnn}. $^3$ \citet{zenkel2018open}. $^4$ \citet{vaswani2018tensor2tensor}. $^5$ \citet{klein2017opennmt}. $^6$ \citet{povey2011kaldi}. $^7$ \citet{pratap2018w2l}.}
    \label{tab:counterparts}
\end{table*}

The increased connections among ASR, ST, MT and LM has called for all-in-one S2S modeling toolkits, and the use of large-scale unlabeled speech data sets the scalability requirements. In this paper, we introduce \stot, a \fairseq~\citep{ott2019fairseq} extension for S2T tasks such as end-to-end ASR and ST. It follows \fairseq's careful design for scalability and extensibility. We provide end-to-end workflows from data pre-processing, model training to offline (online) inference. We implement state-of-the-art RNN-based~\citep{chan2016listen,berard2018end}, Transformer-based~\citep{vaswani2017attention,Mohamed2019TransformersWC} and Conformer-based~\citep{gulati20_interspeech} models and open-source detailed training recipes. \fairseq's MT models and LMs can be seamlessly integrated into S2T workflows for multi-task learning or transfer learning. To facilitate model evaluation, we add a collection of scorers as well as VizSeq~\citep{wang-etal-2019-vizseq} integration for visualized error analysis. \stot~documentation and examples are available at \url{https://github.com/pytorch/fairseq/tree/master/examples/speech_to_text}.

With counterpart toolkits such as ESPNet~\citep{inaguma-etal-2020-espnet} and Lingvo~\citep{shen2019lingvo}, \stot~pursues the best integration, scalability and reproducibility. A detailed comparison of \stot~with its counterparts can be found in Table~\ref{tab:counterparts}.

\section{Features}

\paragraph{Fairseq Models} \fairseq~provides a collection of MT models~\citep{ng-etal-2019-facebook,lewis2019bart} and LMs~\citep{liu2019roberta,conneau2019unsupervised} that demonstrate state-of-the-art performance on standard benchmarks. They are open-sourced with pre-trained models. \fairseq~also supports other tasks such as text summarization, story generation and self-supervised speech pre-training.

\paragraph{S2T extension} \stot~adds attention-based RNN models~\citep{chan2016listen,berard2018end}, Transformer models~\citep{vaswani2017attention,Mohamed2019TransformersWC} as well as the latest Conformer models~\citep{gulati20_interspeech} for ASR and ST. It also supports CTC criterion~\citep{graves2006connectionist} for ASR. For the simultaneous ST setting, it includes online models with widely used policies: monotonic attention~\citep{raffel2017online}, wait-$k$~\citep{Ma2018STACLST},
monotonic infinite lookback attention~\citep{arivazhagan2019monotonic},
and monotonic multihead attention~\citep{ma2019monotonic}.

\paragraph{Data Pre-Processing}
\stot~extracts Kaldi-compliant~\citep{povey2011kaldi} speech features (e.g. log mel-filter banks) automatically from WAV/FLAC audio files via PyKaldi~\citep{pykaldi} or torchaudio\footnote{https://github.com/pytorch/audio}.
Speech features can also be pre-computed and stored in NumPy~\citep{harris2020array} format. Optionally, raw audio files or features files can be packed into ZIP archives to improve I/O performance or facilitate file management. For further pre-processing, \stot~provides online speech data transforms, including CMVN (cepstral mean and variance normalization), speed perturbation~\citep{ko2017study} and SpecAugment~\citep{park2019specaugment}. It also has an open interface for user-defined transforms. For text data, \stot~does online tokenization with a rich collection of tokenizers, including Moses\footnote{https://github.com/moses-smt/mosesdecoder}, SentencePiece~\citep{kudo-richardson-2018-sentencepiece}, subword-nmt\footnote{https://github.com/rsennrich/subword-nmt}, byte-level BPE~\citep{wang2020aaai} and bytes~\citep{li2019bytes}.

\paragraph{Data Configuration} \stot~gets raw audio (feature) paths and target texts from manifest files in TSV (tab-separated values) format, which is similar to Kaldi-style scp files. Online speech data transforms and other data-related settings (e.g. tokenizer type and vocabulary) are defined by a separate configuration file in YAML format.

\paragraph{Computation}
\fairseq~is implemented in PyTorch~\citep{paszke2019pytorch} and it provides efficient batching, mixed precision training~\citep{micikevicius2018mixed}, multi-GPU as well as multi-machine training for computational efficiency on large-scale experiments.

\paragraph{Evaluation Metrics} \stot~provides common automatic metrics for ASR, ST and MT, including WER (word error rate), BLEU~\citep{papineni2002bleu} and chrF~\citep{popovic2015chrf}. It also integrates \textsc{simuleval}~\citep{ma2020simuleval} for simultaneous ST/MT metrics such as AL (average lagging)~\citep{Ma2018STACLST} and DAL (differentiable average Lagging)~\citep{cherry2019thinking}.

\paragraph{Visualization} \fairseq~supports Tensorboard\footnote{https://github.com/tensorflow/tensorboard} for monitoring holistic metrics during model training. It also has VizSeq~\citep{wang-etal-2019-vizseq} integration for sequence-level error analysis, where speech and target/predicted text data are visualized with alignments in Jupyter Notebook interface.

\begin{table*}[t]
    \centering
    \small
    \begin{tabular}{rr|cccccccc}
    \toprule
    & & De & Nl & Es & Fr & It & Pt & Ro & Ru \\
    \midrule
    \multicolumn{2}{r|}{Transformer$^1$} & 17.3 & 18.8 & 20.8 & 26.9 & 16.8 & 20.1 & 16.5 & 10.5 \\
    \multicolumn{2}{r|}{Transformer$^{2\dagger}$} & 22.9 & 27.4 & 28.0 & 32.7 & 23.8 & 28.0 & 21.9 & 15.8  \\
    \midrule
    \multicolumn{2}{r|}{T-Sm} & 22.7 & 27.3 & 27.2 & 32.9 & 22.7 & 28.1 & 21.9 & 15.3 \\
    \multicolumn{2}{r|}{Multi. T-Md$^\ast$} & 24.5 & 28.6 & 28.2 & 34.9 & 24.6 & 31.1 & 23.8 & 16.0 \\
    \midrule
    \multirow{4}{*}{\rotatebox[origin=c]{90}{B-Base}} & Offline & 19.2	&23.5&	24.0&	29.1&	16.4&	23.5&	19.7&	13.7 \\
    & High Lat.$^\ddagger$ & 18.6 (6.8) &  22.9 (6.9) &  22.3 (6.8)& 28.4 (6.7)& 15.4 (6.8)& 22.6 (6.9)&  19.1 (6.7)& 12.9 (6.9)\\
    & Mid Lat.$^\ddagger$ & 14.1 (5.4) &  17.9 (5.4) &  17.2 (5.5)& 25.0 (5.3)& 12.0 (5.5)& 17.7 (5.8)&  15.0 (5.6)& 7.2 (5.8)\\
    & Low Lat.$^\ddagger$
    & 8.2 (2.9)  &  12.3 (2.8) &  13.0 (3.0)& 21.1 (2.8)& 6.7 (2.9)& 13.3 (2.9)&  12.1 (2.9)& 4.9 (2.7)\\
    \bottomrule
    \end{tabular}
    \caption{\stot~models on MuST-C. Test BLEU reported (for online models, AL is shown in parentheses). $^1$~\citet{9004003}. $^2$~\citet{inaguma-etal-2020-espnet}.  $^\dagger$ Applied additional techniques: speed perturbation, pre-trained decoder from MT and auxiliary CTC loss for ASR pre-training. $^\ddagger$ Online models using beam size of 1 (instead of 5). $^\ast$ Trained jointly on all 8 languages.}
    \label{tab:mustc_st_results}
\end{table*}
\begin{table}[t]
    \centering
    \small
    \begin{tabular}{r|ccc}
    \toprule
     &  Type & Config. & Params \\
    \midrule
    B-Base & \multirow{2}{*}{RNN$^1$} & 512d, 3L enc./2L dec. & 31M \\
    B-Big &  & 512d, 5L enc./3L dec. & 52M \\
    \midrule
    T-Sm & Trans- & 256d, 12L enc./6L dec. & 31M \\
    T-Md & former$^2$ & 512d, 12L enc./6L dec. & 72M \\
    T-Lg & & 1024d, 12L enc./6L dec. & 263M \\
    % T-XL & & 1024d, 16L enc./6L dec. & 318M \\
    \midrule
    W-Lg & wav2vec & 1024d, 24L & 315M \\
    CW-Lg & 2.0$^3$ & 1024d, 24L, Conformer$^4$ & 618M \\
    \bottomrule
    \end{tabular}
    \caption{\stot~models for benchmarking. For simplicity, we use the same (default) model hyper-parameters and learning rate schedule across all experiments. $^1$~\citet{berard2018end}.
    $^2$~\citet{vaswani2017attention}. $^3$~\citet{NEURIPS2020_92d1e1eb}. $^4$~\citet{gulati20_interspeech}.
    }
    \label{tab:models}
\end{table}

\begin{table}[t]
    \centering
    \small
    \begin{tabular}{r|cc|cc}
    \toprule
    & \multicolumn{2}{c|}{Dev} & \multicolumn{2}{c}{Test} \\
    & Clean & Other & Clean & Other \\
    \midrule
    \multicolumn{5}{c}{100h labeled} \\
    \midrule
    W-Lg$^3$ & 3.3 & 6.5 & 3.1 & 6.3 \\
    \midrule
    T-Sm & 14.0 & 28.7 & 15.3 & 29.6 \\
    + CTC Aux. & 11.8 & 26.8 & 13.9 & 27.3 \\
    CW-Lg & 2.5 & 5.0 & 2.5 & 5.0 \\
    \midrule
    \multicolumn{5}{c}{960h labeled} \\
    \midrule
    LAS$^1$ & - & - & 2.8 & 6.8 \\
    Transformer$^2$ & 2.5 & 6.7 & 2.9 & 7.0 \\
    W-Lg$^3$ & 2.1 & 4.5 & 2.2 & 4.5 \\
    CW-Lg$^4$ & 1.7 & 3.5 & 1.7 & 3.5 \\
    \midrule
    B-Big & 3.7 & 11.4 & 3.9 & 11.5 \\
    T-Sm & 3.8 & 8.9 & 4.4 & 9.0 \\
    T-Md & 3.2 & 8.0 & 3.4 & 7.9 \\
    T-Lg & 3.0 & 7.5 & 3.2 & 7.5 \\
    CW-Lg & 1.7 & 3.5 & 1.8 & 3.7 \\
    \bottomrule
    \end{tabular}
    \caption{\stot~models on LibriSpeech. Dev and test WER reported. $^1$~\citet{park2019specaugment}. $^2$~\citet{synnaeve2019end}. $^3$~\citet{NEURIPS2020_92d1e1eb}. $^4$~\citet{zhang2020pushing}.}
    \label{tab:librispeech_results}
\end{table}

\begin{table*}[t]
    \small
    \centering
    \begin{tabular}{r|ccccccccccccc}
    \toprule
         & Fr & De & Es & Zh & Tr & Ar & Sv & Lv & Sl & Ta & Ja & Id & Cy \\
         \midrule
         \multicolumn{14}{c}{X$\rightarrow$En} \\
         \midrule
         B-Base & 23.2 & 15.7 & 20.2 & 4.4 & 2.2 & 2.7 & 1.4 & 1.2 & 1.5 & 0.2 & 1.1 & 1.0 & 1.7 \\
         + SSL$^\star$ & 23.1& 16.2& 20.2& 4.8& 3.2& 3.8& 3.7& 2.3& 2.2& 0.2& 1.6& 1.6& 2.2 \\
         Multi. B-Big$^\dagger$ & 26.6 & 19.5 & 26.3 & 4.4 & 2.1 & 0.3 & 1.3 & 0.6 & 1.4 & 0.1 & 0.6 & 0.3 & 0.9 \\
         T-Sm & 26.3 & 17.1 & 23.0 & 5.8 & 3.6 & 4.3 & 2.7 & 2.5 & 3.0 & 0.3 & 1.5 & 2.5 & 2.7 \\
         Multi. T-Md$^\dagger$ & 26.5 & 17.5 & 27.0 & 5.9 & 2.3 & 0.4 & 0.5 & 0.6 & 0.7 & 0.1 & 0.1 & 0.3 & 1.9 \\
         \midrule
         \multicolumn{14}{c}{En$\rightarrow$X} \\
         \midrule
         B-Base & - & 12.5 & - & 20.0 & 6.7 & 9.1 & 18.1 & 8.7 & 11.6 & 7.4 & 25.6 & 15.2 & 18.9 \\
         Multi. B-Big$^\ddagger$ & - & 12.6 & - & 22.2 & 7.3 & 8.0 & 18.3 & 8.9 & 11.4 & 7.3 & 28.2 & 16.0 & 19.3 \\
         T-Sm & - & 16.3 & - & 25.4 & 10.0 & 12.1 & 21.8 & 13.0 & 16.0 & 10.9 & 29.6 & 20.4 & 23.9 \\
         Multi. T-Md$^\ddagger$ & - & 15.4 & - & 26.5 & 9.5 & 10.8 & 20.9 & 12.2 & 14.6 & 10.3 & 30.5 & 18.9 & 22.0 \\
    \bottomrule
    \end{tabular}

    \caption{\stot~models on CoVoST 2. Test BLEU reported (character-level BLEU for Zh and Ja targets). $^\star$ Replaced mel-filter bank features with wav2vec ones ~\citep{Schneider2019,wu2020selfsupervised}. $^\dagger$ Trained jointly on all 21 X-En directions with temperature-based (T=2) resampling~\citep{arivazhagan2019massively}. $^\ddagger$ Trained jointly on all 15 En-X directions.}
    \label{tab:covost2_results}
\end{table*}

\section{Experiments}
We evaluate \stot~models on English ASR benchmark---LibriSpeech~\citep{panayotov2015librispeech}, as well as multilingual ST benchmarks---MuST-C~\citep{di2019must} and CoVoST 2~\citep{wang2020covost}. The model architectures used in benchmarking can be found in Table~\ref{tab:models}.

\subsection{Experimental Setup}
For speech inputs, we extract 80-channel log mel-filter bank features (25ms window size and 10ms shift) with utterance-level CMVN applied. We remove training samples with more than 3,000 frames for GPU memory efficiency. To alleviate overfitting, we pre-train ST model encoders on English ASR and adopt SpecAugment (without time warping): LD policy on LibriSpeech models and LB policy on MuST-C and CoVoST 2 models. We average the last 10 checkpoints and use a beam size of 5 for decoding. For ASR, we use 10K unigram vocabulary~\citep{kudo-richardson-2018-sentencepiece} and report WER. For ST, we use character vocabulary for CoVoST 2 and 8K unigram vocabulary for MuST-C. We report case-sensitive detokenized BLEU using sacreBLEU~\cite{post-2018-call}, except for Japanese and Chinese translations (no word segmentation) where we report character-level BLEU.

\subsection{Speech Recognition (ASR)}
LibriSpeech is a de-facto standard ASR benchmark that contains 1,000 hours of English speech from audiobooks. Table~\ref{tab:librispeech_results} shows the dev and test WER of our models on LibriSpeech clean and noisy sets. Three architectures, RNN-based model (``B-Big"), Transformer-based models (``T-Sm", ``T-Md" and ``T-Lg") and Conformer-based wav2vec 2.0 model (``CW-Lg"), are evaluated. We can see that the first two architectures are able to achieve competitive performance (WER) to the state-of-the-art ones, while we use only default model hyper-parameters and learning rate schedule without any task-specific tuning. Our implementation of the third architecture matches the state of the art.

\subsection{Speech Translation (ST)}

\subsubsection{MuST-C}
MuST-C contains up to around 500 hours of English speech from TED talks with translations in 8 European languages.
Table~\ref{tab:mustc_st_results} shows the test BLEU of our Transformer-based models (``T-Sm" and ``Multi. T-Md") and RNN-based models (``B-Base") on all the MuST-C language directions.
Compared with previous Transformer-based approaches \cite{di-gangi-etal-2019-enhancing, inaguma-etal-2020-espnet}, our bilingual models achieve comparative results to the state of the art without applying additional techniques such as speed perturbation and pre-trained decoder from MT. Moreover, our multilingual model (trained on all 8 languages) outperforms all bilingual ones with large margins. Besides traditional offline models, we also provide simultaneous ST models: the lower section in Table~\ref{tab:mustc_st_results} presents the online models with wait-$k$ policy, which was the baseline system in the IWSLT 2020 shared task on simultaneous ST~\citep{ansari-etal-2020-findings}. 
The results represent the best systems in
high ($\textrm{AL} > 6$), medium ($6 \ge \textrm{AL} > 3$) and low ($\textrm{AL} \le 3$) latency regimes, on which we can clearly see the trade-offs between model performance and prediction latency.

\subsubsection{CoVoST 2}
CoVoST 2 contains total 2,880 hours of read speech in 22 languages from the open-source community, with 21 X-En directions and 15 En-X directions. We evaluate our models bidirectionally on 13 languages of them, including low-resource X-En directions: Zh, Tr, Ar, Sv, Lv, Sl, Ta, Ja, Id and Cy. We observe from Table~\ref{tab:covost2_results} that our Transformer-based models (``T-Sm" and ``T-Md") outperforms RNN-based ones (``B-Base" and ``B-Big") on all En-X and X-En directions. The performance gap tends to be larger when the training data is higher resource (En-X directions, Fr-En, De-En and Es-En). Our multilingual models perform reasonably well with a universal model for over 15 X-En or En-X directions. They even have significant improvements on some directions (e.g. at least 4 BLEU gain on Es-En). For low-resource directions, we also evaluate self-supervised speech features~\citep{Schneider2019,wu2020selfsupervised}\footnote{From a wav2vec model pre-trained on LibriSpeech: \url{https://github.com/pytorch/fairseq/tree/master/examples/wav2vec}} as an alternative to the traditional log mel-filter bank features (``+ SSL"). We find that self-supervised features bring consistent gains and transfer well across different languages (self-supervised model trained on English and feature extracted for non-English).

\section{Conclusion}
We introduce \stot, a \fairseq~extension for speech-to-text (S2T) modeling tasks such as speech recognition and speech translation. It includes end-to-end workflows and state-of-the-art models with scalablity and extensibility design. It seamlessly integrates \fairseq's machine translation models and language models to improve S2T model performance. \stot~documentation and examples are available at \url{https://github.com/pytorch/fairseq/tree/master/examples/speech_to_text}.

\section*{Acknowledgments}
We thank Myle Ott, Michael Auli, Alexei Baevski, Jiatao Gu, Abdelrahman Mohamed and Javad Dousti for helpful discussions.

\bibliography{aacl-ijcnlp2020}
\bibliographystyle{acl_natbib}

\end{document}